How much room is there to spread frosting on the cookie? Clare says the radius of the cookie is about 3 cm so the space for frosting is about 6n cm Andre says the diameter of the cookie is about 3 inches so the space for frosting is about 2.25in is this question talking about area or circumference?
1
Expert's answer
2020-04-30T20:01:42-0400

This question is talking about area.


"Area=\pi r^2={\pi d^2 \over 4}"

If the radius of the cookie is about 3 cm, the space for frosting is about "9\pi" cm2. If the diameter of the cookie is about 3 inches the space for frosting is about "2.25\pi" in2.